OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 3pi

OpenStudy (anonymous):

Aaaaaaaargh!

OpenStudy (anonymous):

Oops, typed it wrong 2*pi*r

OpenStudy (anonymous):

i trieddoing it myself i have i did 2*3.14*3 and it gaint 18.84

OpenStudy (anonymous):

Third time lucky...."exact" means leave pi as pi.

OpenStudy (anonymous):

to get the radius, 2*pi*r=3pi r =(3pi)/(2pi) {cancel pi and pi) so r = 3/2 area=pi*r^2 area=(22/7)*(3/2)*(3/2) area=(11/7)*3*(3/2) area=99/14 area = 7.07